# Unifying Appearance Codes and Bilateral Grids for Driving Scene Gaussian Splatting

### v1

@@ Abstract
Neural rendering techniques, including NeRF and Gaussian Splatting (GS), rely on photometric consistency to produce high-quality reconstructions. However, in real-world scenarios, it is challenging to guarantee perfect photometric consistency in acquired images. Appearance codes have been widely used to address this issue, but their modeling capability is limited, as a single code is applied to the entire image. Recently, the bilateral grid was introduced to perform pixel-wise color mapping, but it is difficult to optimize and constrain effectively. In this paper, we propose a novel multi-scale bilateral grid that unifies appearance codes and bilateral grids. We demonstrate that this approach significantly improves geometric accuracy in dynamic, decoupled autonomous driving scene reconstruction, outperforming both appearance codes and bilateral grids. This is crucial for autonomous driving, where accurate geometry is important for obstacle avoidance and control. Our method shows strong results across four datasets: Waymo, NuScenes, Argoverse, and PandaSet. We further demonstrate that the improvement in geometry is driven by the multi-scale bilateral grid, which effectively reduces floaters caused by photometric inconsistency. Our code is open-sourced at https://bigcileng.github.io/bilateral-driving.
Figure 1: (a) Appearance codes rely on global affine transformations but have limited modeling capability. (b) Bilateral grids perform pixel-wise transformations, improving photometric consistency but are challenging to optimize. (c) The proposed multi-scale bilateral grid unifies appearance codes and bilateral grids, enabling patch-wise transformations.
@@ 1 Introduction
Neural rendering techniques, such as NeRFs  mildenhall2021nerf ; wang2021neus ; yu2022monosdf ; liu2024rip ; yuan2024slimmerf and Gaussian Splatting (GS) kerbl20233d ; chen2024pgsr ; cheng2024gaussianpro ; zhang2024drone ; song2024sa , have demonstrated significant potential in producing high-quality 3D reconstructions by leveraging photometric consistency. However, ensuring photometric consistency across images in real-world scenarios remains a challenge, as lighting conditions, viewpoints, and camera settings can vary considerably martin2021nerf ; zhang2024gaussian ; kulhanek2024wildgaussians ; wang2025semantic . In autonomous driving applications yan2024renderworld ; wu2023mars ; tonderski2024neurad ; chen2024omnire ; li2024uniscene ; jin2024tod3cap ; tian2023unsupervised , these challenges are amplified due to the presence of multiple cameras capturing the same scene from different angles over time. As illustrated in the left half of Teaser. 1, we observe that these variations—whether temporal (across different time steps) or inter-camera (across different viewpoints)—can introduce significant discrepancies in image appearance. These inconsistencies pose difficulties in accurately reconstructing dynamic driving environments.
@@
To address this issue, appearance codes tonderski2024neurad ; martin2021nerf ; chen2024omnire have been introduced to encode per-image information and assist in correcting the photometric discrepancies. These codes, while effective, have limitations in modeling capability since they apply a single transformation across the entire image, often neglecting local variations that may occur in complex scenes. For example, in dynamic scenes, such as those encountered in autonomous driving environments, lighting and viewpoints can change across different camera angles and time frames. As shown in the right top of Teaser. 1, a global affine transformation based on a single appearance code may not be sufficient to handle such variations (In Teaser. 1 NuScenes (a), the fence is blurred).
@@
Recent advancements wang2024bilateral ; gharbi2017deep have introduced bilateral grids to perform pixel-wise transformations, enabling improved photometric consistency by allowing more localized adjustments. However, these grids face significant challenges in optimization, as they require complex constraints to avoid instability during training (As shown in Teaser. 1 (b), bilateral grids fail to effectively optimize the reconstruction of large, complex scenes.). In this work, we propose a novel solution—a multi-scale bilateral grid—integrating the strengths of appearance codes and bilateral grids. As shown in Teaser. 1, this new approach allows for patch-wise transformations. Interestingly, in extreme cases, the multi-scale bilateral grid naturally converges to either the bilateral grid or appearance code, depending on the scale. At the finest scale, the multi-scale grid behaves like a traditional bilateral grid, performing pixel-wise transformations to adjust for local variations. In contrast, at the coarsest scale, it effectively reverts to a more global transformation, resembling the behavior of appearance codes. This flexible, scale-dependent approach offers the best of both worlds, providing localized fine-tuning where necessary, while maintaining global consistency when appropriate.
@@
We demonstrate that the multi-scale bilateral grid significantly enhances the geometric accuracy of dynamic, decoupled driving scenes, which is essential for autonomous driving applications where precise geometry plays a crucial role in tasks like obstacle avoidance and path planning. This is evidenced by our evaluation on multiple widely used datasets, including Waymo sun2020scalability , NuScenes caesar2020nuscenes , Argoverse wilson2023argoverse , and PandaSet xiao2021pandaset . The corresponding Chamfer Distance (CD) values, displayed in the Teaser. 1, highlight a marked reduction in error and improvement in reconstruction quality. Additionally, our method reduces photometric inconsistency, mitigating the appearance of floaters and enhancing the overall realism of the scene.
@@
Figure 2: Overview of our method. We unify appearance codes with multi-scale bilateral grids. Initially, a coarse rendering is obtained from a Gaussian scene graph. This rendered image is then processed by our multi-scale bilateral grids to perform detailed per-pixel color modeling, guided by a luminance-based map through slice and fusion operations.
@@ 2.1 Bilateral Grids and Appearance Codes
Recent advancements in bilateral grids and appearance codes have significantly influenced techniques for addressing photometric inconsistencies, which are particularly critical in dynamic environments such as autonomous driving. The bilateral grids, first introduced for real-time edge-aware image processing chen2007real , enable efficient manipulation of spatial and intensity variations, forming the foundation for many modern techniques gharbi2017deep ; marki2016bilateral ; chen2016bilateral ; xu2021bilateral ; zheng2021ultra . Extensions like deep bilateral learning gharbi2017deep further enhance this framework by predicting local affine transformations in bilateral space, achieving real-time image enhancement on resource-constrained devices. Similarly, bilateral-guided upsampling chen2016bilateral and cost volume refinement xu2021bilateral have proven effective in tasks ranging from high-dynamic-range imaging to stereo matching, demonstrating the versatility of bilateral representations. In the context of neural rendering, methods such as NeRF tancik2022block ; jang2021codenerf ; martin2021nerf and Gaussian Splatting (GS) zhang2024gaussian ; kulhanek2024wildgaussians have achieved groundbreaking results in novel view synthesis but photometric inconsistency caused by varying illumination or transient occluders martin2021nerf ; zhang2024gaussian remains a challenge. Techniques like bilateral guided radiance field processing wang2024bilateral address these issues by disentangling camera-specific enhancements and reapplying them consistently in 3D space. WildGaussians kulhanek2024wildgaussians integrates DINO-based appearance codes into 3D GS to robustly handle occlusions and dynamic lighting in uncontrolled scenes. Cross-Ray NeRF yang2023cross employs cross-ray feature covariance and grid-based transient masking to harmonize appearance variations and suppress occlusions in unconstrained image collections. Despite these strides, existing approaches often struggle to balance global consistency with localized adaptability. For instance, while appearance codes provide global adjustments, they lack the granularity to model fine-grained variations durand2002fast . Conversely, bilateral grids excel at pixel-wise transformations but are challenging to optimize effectively paris2006fast . To address this issue, we introduce a multi-scale bilateral grid which unifies bilateral grids and appearance codes, facilitating high-quality dynamic autonomous driving scene reconstruction with enhanced modeling capability and optimization efficiency.
@@ 2.2 Autonomous Driving Simulation
Autonomous driving simulation has emerged as a critical tool for developing perception, planning, and control systems by generating diverse, realistic driving scenarios wu2023mars ; okae2021robust ; yan2024street ; lindstrom2024nerfs ; tao2024alignmif ; li2024uniscene . Recent advancements focus on photorealistic rendering, dynamic scene modeling, and multi-modal sensor simulation wu2023mars ; li2024hierarchical ; zhou2024drivinggaussian ; tonderski2024neurad ; yan2024street . For instance, MARS wu2023mars employs a modular NeRF-based framework to independently control static and dynamic scene elements, while Street Gaussian yan2024street achieves real-time urban scene rendering using explicit 3D Gaussian representations. DrivingGaussian zhou2024drivinggaussian enhances dynamic scene reconstruction via composite Gaussian Splatting, ensuring occlusion accuracy and multi-camera consistency. NeuRAD tonderski2024neurad integrates sensor-specific effects (e.g., rolling shutter, LiDAR beam divergence) to improve novel view synthesis. Holistic scene understanding and editing have also seen progress. HUGS zhou2024hugs combines static and dynamic 3D Gaussians for real-time semantic parsing. ChatSim wei2024editable enables language-driven scene editing with external asset integration. LiDAR data integration has advanced through methods like LiDAR-NeRF tao2024lidar , which uses structural regularization for low-texture regions. HO-Gaussian li2024ho , which merges grid-based volumes with Gaussian Splatting to eliminate Structure-from-Motion (SFM) dependencies.
Despite progress, gaps persist between simulated and real-world data lindstrom2024nerfs ; tao2024alignmif . AlignMiF tao2024alignmif addresses LiDAR-camera misalignments via geometry-aligned implicit fields, while RodUS nguyen2024rodus decomposes urban scenes into static and dynamic components using 4D semantics to reduce artifacts. Innovations in large-scale generation include InfiniCube lu2024infinicube , which leverages sparse-voxel representations for unbounded dynamic scenes, and Omnire chen2024omnire , which reconstructs diverse dynamic objects (e.g., pedestrians) for human-vehicle interaction simulations. Furthermore, recent works like GaussianPro cheng2024gaussianpro and SplatAD hess2024splatad explore to refine 3D GS for real-time rendering of dynamic scenes. GaussianPro cheng2024gaussianpro introduces progressive propagation for texture-less surfaces, and SplatAD hess2024splatad models sensor-specific phenomena (e.g., rolling shutter, LiDAR intensity). In this work, we aim to reconstruct high-quality driving scenes by unifying appearance codes and bilateral grids to enhance photometric consistency and modeling capability.
@@ 3.2 Gaussian Splatting for Autonomous Driving Environments
Following prior methodologies chen2024omnire ; tonderski2024neurad , we represent autonomous driving environments using a hybrid scene graph, decomposing the scene into sky, background, and dynamic object models. Each dynamic object is represented by a 3D model in canonical space and transformed to the scene with an associated sequence of SE(3) transformations. We derive the transformation matrixes from existing object detection pipelines or ground-truth annotations.
@@ 3.3.2 Guidance Map, Slice Operation, and Multi-Scale Fusion
To achieve adaptive and spatially-varying photometric correction, our multi-scale bilateral grid framework employs a guidance map and a slicing operation chen2007real to retrieve localized transformations, followed by a hierarchical fusion strategy to combine transformations across scales.
We first derive a luminance-based guidance map $I^{g}(u,v)=\text{GrayScale}(I^{r}(u,v))$ from the rendered image $I^{r}(u,v)$, following the approach of chen2007real ; wang2024bilateral ; gharbi2017deep . This map encodes spatial brightness variations like shadows and highlights, serving as a spatially-varying query to our multi-scale grid. For each level $l$ and pixel $(u,v)$, the luminance $d$ from $I^{g}$ is used to perform a slicing operation, querying the grid’s intensity dimension $D^{(l)}$ to locate an intensity bin. Affine transformation coefficients $\mathcal{A}^{(l)}_{i,j,k}$ around this bin are then combined via trilinear interpolation for a level-specific transformation $\bar{A}^{(l)}(u,v)$. The slicing operation retrieves per-pixel transformations from each grid level, written as:
$\bar{A}^{(l)}(u,v)=\sum\limits_{i,j,k}w_{i,j,k}(u,v,d)\mathcal{A}^{s}(i,j,k)~{}~{},$ (7)
@@
Dataset Method Full Image Human Vehicle Geometry
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ CD $\downarrow$ RMSE $\downarrow$ Depth $\downarrow$
Waymo sun2020scalability Baseline 28.92 0.833 25.24 0.681 23.50 0.697 1.482 2.785 0.540
Baseline w/AC 28.95 0.835 24.92 0.656 23.51 0.699 1.378 2.760 0.519
Baseline w/BG 28.19 0.831 24.79 0.664 23.09 0.693 1.523 2.798 0.492
Ours 29.23 0.836 25.55 0.683 23.50 0.707 0.989 2.744 0.477
NuScenes caesar2020nuscenes Baseline 26.37 0.837 26.17 0.774 26.49 0.827 1.458 3.420 0.110
Baseline w/AC 26.38 0.840 25.83 0.762 26.52 0.828 1.437 3.415 0.106
Baseline w/BG 25.98 0.837 26.00 0.771 26.10 0.820 1.380 3.390 0.097
Ours 27.69 0.847 26.53 0.777 27.31 0.834 1.161 3.340 0.059
Argoverse wilson2023argoverse Baseline 24.59 0.848 23.35 0.743 24.47 0.812 0.954 4.208 0.050
Baseline w/AC 24.58 0.848 23.22 0.738 24.44 0.811 0.959 4.215 0.051
Baseline w/BG 23.31 0.842 22.99 0.730 23.49 0.795 0.901 4.218 0.049
Ours 24.68 0.849 23.35 0.739 24.51 0.827 0.807 4.199 0.040
Pandaset xiao2021pandaset Baseline 30.20 0.903 29.79 0.830 23.27 0.776 0.503 2.874 0.018
Baseline w/AC 30.20 0.903 29.64 0.824 23.39 0.783 0.496 2.871 0.020
Baseline w/BG 29.73 0.904 29.90 0.812 22.22 0.763 0.484 2.867 0.013
Ours 30.75 0.906 30.18 0.830 23.49 0.784 0.453 2.852 0.011
Table 1: Comparison with SOTA for scene reconstruction. w/AC represents appearance codes while w/BG represents single bilateral grids.

### v4

@@ Abstract
Neural rendering techniques, including NeRF and Gaussian Splatting (GS), rely on photometric consistency to produce high-quality reconstructions. However, in real-world driving scenarios, it is challenging to guarantee perfect photometric consistency in acquired images. Appearance codes have been widely used to address this issue, but their modeling capability is limited, as a single code is applied to the entire image. Recently, the bilateral grid was introduced to perform pixel-wise color mapping, but it is difficult to optimize and constrain effectively. In this paper, we propose a novel multi-scale bilateral grid that unifies appearance codes and bilateral grids. We demonstrate that this approach significantly improves geometric accuracy in dynamic, decoupled autonomous driving scene reconstruction, outperforming both appearance codes and bilateral grids. This is crucial for autonomous driving, where accurate geometry is important for obstacle avoidance and control. Our method shows strong results across four datasets: Waymo, NuScenes, Argoverse, and PandaSet. We further demonstrate that the improvement in geometry is driven by the multi-scale bilateral grid, which effectively reduces floaters caused by photometric inconsistency. Our code is open-sourced at https://bigcileng.github.io/bilateral-driving.
Figure 1: Unifying appearance codes and bilateral grids. (a) Appearance codes rely on global affine transformations but have limited modeling capability. (b) Bilateral grids perform pixel-wise transformations, improving photometric consistency but are challenging to optimize. (c) The proposed multi-scale bilateral grid unifies both paradigms, enabling patch-wise transformations.
@@ 1 Introduction
Neural rendering techniques, such as NeRFs  Mildenhall et al. (2021); Wang et al. (2021); Yu et al. (2022); Liu et al. (2024); Yuan and Zhao (2024) and Gaussian Splatting (GS) Kerbl et al. (2023); Chen et al. (2024a); Cheng et al. (2024); Zhang et al. (2024b); Song et al. (2024); Ye et al. (2025); Shuai et al. (2025), have demonstrated significant potential in producing high-quality 3D reconstructions by leveraging photometric consistency. However, ensuring photometric consistency across images in real-world scenarios remains a challenge, as lighting conditions, viewpoints, and camera settings can vary considerably Martin-Brualla et al. (2021); Zhang et al. (2024a); Kulhanek et al. (2024); Wang et al. (2025); Gao et al. (2024). In autonomous driving applications Yan et al. (2024b); Wu et al. (2023); Tonderski et al. (2024); Chen et al. (2024b); Li et al. (2024b); Jin et al. (2024); Tian et al. (2023); Han et al. (2025); Yu et al. (2025), these challenges are amplified due to the presence of multiple cameras capturing the same scene from different angles over time. As illustrated in the left half of Teaser. 1, we observe that these variations—whether temporal (across different time steps) or inter-camera (across different viewpoints)—can introduce significant discrepancies in image appearance. These inconsistencies pose difficulties in accurately reconstructing dynamic driving environments.
@@
To address this issue, appearance codes Tonderski et al. (2024); Martin-Brualla et al. (2021); Chen et al. (2024b) have been introduced to encode per-image information and assist in correcting the photometric discrepancies. These codes, while effective, have limitations in modeling capability since they apply a single transformation across the entire image, often neglecting local variations that may occur in complex scenes. For example, in dynamic scenes, such as those encountered in autonomous driving environments, lighting and viewpoints can change across different camera angles and time frames. As shown in the right top of Teaser. 1, a global affine transformation based on a single appearance code may not be sufficient to handle such variations (In Teaser. 1 NuScenes (a), the fence is blurred).
@@
Recent advancements Wang et al. (2024); Gharbi et al. (2017) have introduced bilateral grids to perform pixel-wise transformations, enabling improved photometric consistency by allowing more localized adjustments. However, these grids face significant challenges in optimization, as they require complex constraints to avoid instability during training (As shown in Teaser. 1 (b), bilateral grids fail to effectively optimize the reconstruction of large, complex scenes.). In this work, we propose a novel solution—a multi-scale bilateral grid—integrating the strengths of appearance codes and bilateral grids. As shown in Teaser. 1, this new approach allows for patch-wise transformations. Interestingly, in extreme cases, the multi-scale bilateral grid naturally converges to either the bilateral grid or appearance code, depending on the scale. At the finest scale, the multi-scale grid behaves like a traditional bilateral grid, performing pixel-wise transformations to adjust for local variations. In contrast, at the coarsest scale, it effectively reverts to a more global transformation, resembling the behavior of appearance codes. This flexible, scale-dependent approach offers the best of both worlds, providing localized fine-tuning where necessary, while maintaining global consistency when appropriate.
@@
We demonstrate that the multi-scale bilateral grid significantly enhances the geometric accuracy of dynamic, decoupled driving scenes, which is essential for autonomous driving applications where precise geometry plays a crucial role in tasks like obstacle avoidance and path planning. This is evidenced by our evaluation on multiple widely used datasets, including Waymo Sun et al. (2020), NuScenes Caesar et al. (2020), Argoverse Wilson et al. (2023), and PandaSet Xiao et al. (2021). The corresponding Chamfer Distance (CD) values, displayed in the Teaser. 1, highlight a marked reduction in error and improvement in reconstruction quality. Additionally, our method reduces photometric inconsistency, mitigating the appearance of floaters and enhancing the overall realism of the scene.
@@
Figure 2: Overview of our method. We unify appearance codes with multi-scale bilateral grids. A coarse rendering from the Gaussian scene graph is refined by multi-scale bilateral grids to model per-pixel color with a luminance-guided slice-and-fuse pipeline.
@@ 2.1 Bilateral Grids and Appearance Codes
Recent advancements in bilateral grids and appearance codes have significantly influenced techniques for addressing photometric inconsistencies, which are particularly critical in dynamic environments such as autonomous driving. The bilateral grids, first introduced for real-time edge-aware image processing Chen et al. (2007), enable efficient manipulation of spatial and intensity variations, forming the foundation for many modern techniques Gharbi et al. (2017); Märki et al. (2016); Chen et al. (2016); Xu et al. (2021); Zheng et al. (2021). Extensions like deep bilateral learning Gharbi et al. (2017) further enhance this framework by predicting local affine transformations in bilateral space, achieving real-time image enhancement on resource-constrained devices. Similarly, bilateral-guided upsampling Chen et al. (2016) and cost volume refinement Xu et al. (2021) have proven effective in tasks ranging from high-dynamic-range imaging to stereo matching, demonstrating the versatility of bilateral representations. In the context of neural rendering, methods such as NeRF Tancik et al. (2022); Jang and Agapito (2021); Martin-Brualla et al. (2021) and Gaussian Splatting (GS) Zhang et al. (2024a); Kulhanek et al. (2024) have achieved groundbreaking results in novel view synthesis but photometric inconsistency caused by varying illumination or transient occluders Martin-Brualla et al. (2021); Zhang et al. (2024a) remains a challenge. Techniques like bilateral guided radiance field processing Wang et al. (2024) address these issues by disentangling camera-specific enhancements and reapplying them consistently in 3D space. WildGaussians Kulhanek et al. (2024) integrates DINO-based appearance codes into 3D GS to robustly handle occlusions and dynamic lighting in uncontrolled scenes. Cross-Ray NeRF Yang et al. (2023b) employs cross-ray feature covariance and grid-based transient masking to harmonize appearance variations and suppress occlusions in unconstrained image collections. Despite these strides, existing approaches often struggle to balance global consistency with localized adaptability. For instance, while appearance codes provide global adjustments, they lack the granularity to model fine-grained variations Durand and Dorsey (2002). Conversely, bilateral grids excel at pixel-wise transformations but are challenging to optimize effectively Paris and Durand (2006). To address this issue, we introduce a multi-scale bilateral grid which unifies bilateral grids and appearance codes, facilitating high-quality dynamic autonomous driving scene reconstruction with enhanced modeling capability and optimization efficiency.
@@ 2.2 Autonomous Driving Simulation
Autonomous driving simulation has emerged as a critical tool for developing perception, planning, and control systems by generating diverse, realistic driving scenarios Wu et al. (2023); Okae et al. (2021); Yan et al. (2024a); Lindström et al. (2024); Tao et al. (2024b); Li et al. (2024b); Yang et al. (2024a); Gao et al. (2023). Recent advancements focus on photorealistic rendering, dynamic scene modeling, and multi-modal sensor simulation Wu et al. (2023); Li et al. (2024a); Zhou et al. (2024b); Tonderski et al. (2024); Yan et al. (2024a); Yang et al. (2023a, 2025b); Xiao et al. (2025). For instance, MARS Wu et al. (2023) employs a modular NeRF-based framework to independently control static and dynamic scene elements, while Street Gaussian Yan et al. (2024a) achieves real-time urban scene rendering using explicit 3D Gaussian representations. DrivingGaussian Zhou et al. (2024b) enhances dynamic scene reconstruction via composite Gaussian Splatting, ensuring occlusion accuracy and multi-camera consistency. NeuRAD Tonderski et al. (2024) integrates sensor-specific effects (e.g., rolling shutter, LiDAR beam divergence) to improve novel view synthesis. Holistic scene understanding and editing have also seen progress. HUGS Zhou et al. (2024a) combines static and dynamic 3D Gaussians for real-time semantic parsing. ChatSim Wei et al. (2024) enables language-driven scene editing with external asset integration. LiDAR data integration has advanced through methods like LiDAR-NeRF Tao et al. (2024a), which uses structural regularization for low-texture regions. HO-Gaussian Li et al. (2024c) merges grid-based volumes with Gaussian Splatting to eliminate Structure-from-Motion dependencies.
Despite progress, gaps persist between simulated and real-world data Lindström et al. (2024); Tao et al. (2024b); Yang et al. (2024b, 2025a). AlignMiF Tao et al. (2024b) addresses LiDAR-camera misalignments via geometry-aligned implicit fields, while RodUS Nguyen et al. (2024) decomposes urban scenes into static and dynamic components using 4D semantics to reduce artifacts. Innovations in large-scale generation include InfiniCube Lu et al. (2024), which leverages sparse-voxel representations for unbounded dynamic scenes, and Omnire Chen et al. (2024b), which reconstructs diverse dynamic objects (e.g., pedestrians) for human-vehicle interaction simulations. Furthermore, recent works like GaussianPro Cheng et al. (2024) and SplatAD Hess et al. (2024) explore to refine 3D GS for real-time rendering of dynamic scenes. GaussianPro Cheng et al. (2024) introduces progressive propagation for texture-less surfaces, and SplatAD Hess et al. (2024) models sensor-specific phenomena (e.g., rolling shutter, LiDAR intensity). In this work, we aim to reconstruct high-quality driving scenes by unifying appearance codes and bilateral grids to enhance photometric consistency and modeling capability.
@@ 3.2 Gaussian Splatting for Autonomous Driving Environments
Following prior methodologies Chen et al. (2024b); Tonderski et al. (2024), we represent autonomous driving environments using a hybrid scene graph, decomposing the scene into sky, background, and dynamic object models. Each dynamic object is represented by a 3D model in canonical space and transformed to the scene with an associated sequence of SE(3) transformations. We derive the transformation matrixes from existing object detection pipelines or ground-truth annotations.
@@ 3.3.2 Guidance Map, Slice Operation, and Multi-Scale Fusion
To achieve adaptive and spatially-varying photometric correction, our multi-scale bilateral grid framework employs a guidance map and a slicing operation Chen et al. (2007) to retrieve localized transformations, followed by a hierarchical fusion strategy to combine transformations across scales.
We first derive a luminance-based guidance map $I^{g}(u,v)=\text{GrayScale}(I^{r}(u,v))$ from the rendered image $I^{r}(u,v)$, following the approach of Chen et al. (2007); Wang et al. (2024); Gharbi et al. (2017). This map encodes spatial brightness variations like shadows and highlights, serving as a spatially-varying query to our multi-scale grid. For each level $l$ and pixel $(u,v)$, the luminance $d$ from $I^{g}$ is used to perform a slicing operation, querying the grid’s intensity dimension $D^{(l)}$ to locate an intensity bin. Affine transformation coefficients $\mathcal{A}^{(l)}_{i,j,k}$ around this bin are then combined via trilinear interpolation for a level-specific transformation $\bar{A}^{(l)}(u,v)$. The slicing operation retrieves per-pixel transformations from each grid level, written as:
$\bar{A}^{(l)}(u,v)=\sum\limits_{i,j,k}w_{i,j,k}(u,v,d)\mathcal{A}^{(l)}(i,j,k)~~,$ (7)
@@
Table 1: Ablation studies and comparisons on four large-scale driving datasets. Our proposed method consistently outperforms OmniRe across all metrics and datasets. w/AC denotes appearance codes; w/BG denotes single bilateral grids (size of $16\times 16\times 8$).
Dataset Method Reconstruction Novel View Synthesis Geometry
Waymo OmniRe $28.92$ $0.833$ 0.295 $26.40$ 0.761 0.311 $1.482$ $2.785$ $0.540$
OmniRe w/AC $28.95$ $0.835$ 0.293 $26.44$ 0.759 0.315 $1.378$ $2.760$ $0.519$
OmniRe w/BG $28.19$ $0.831$ 0.292 $21.51$ 0.743 0.333 $1.523$ $2.798$ $0.492$
Ours 29.23 0.836 0.289 26.55 0.762 0.310 0.989 2.744 0.477
NuScenes OmniRe $26.37$ $0.837$ 0.209 $23.74$ 0.733 0.232 $1.458$ $3.420$ $0.110$
OmniRe w/AC $26.38$ $0.840$ 0.204 $23.74$ 0.732 0.229 $1.437$ $3.415$ $0.106$
OmniRe w/BG $25.98$ $0.837$ 0.209 $23.60$ 0.705 0.262 $1.380$ $3.390$ $0.097$
Ours 27.69 0.847 0.193 24.64 0.739 0.216 1.161 3.340 0.059
Argoverse OmniRe $24.59$ $0.848$ 0.202 $22.53$ 0.755 0.220 $0.954$ $4.208$ $0.050$
OmniRe w/AC $24.58$ $0.848$ 0.201 $22.51$ 0.756 0.219 $0.959$ $4.215$ $0.051$
OmniRe w/BG $23.31$ $0.842$ 0.216 $21.70$ 0.725 0.254 $0.901$ $4.218$ $0.049$
Ours 24.68 0.849 0.200 22.58 0.756 0.217 0.807 4.199 0.040
PandaSet OmniRe $30.20$ $0.903$ 0.219 $27.49$ 0.835 0.240 $0.503$ $2.874$ $0.018$
OmniRe w/AC $30.20$ $0.903$ 0.220 $27.51$ 0.841 0.242 $0.496$ $2.871$ $0.020$
OmniRe w/BG $29.73$ $0.904$ 0.220 $27.38$ 0.830 0.246 $0.484$ $2.867$ $0.013$
Ours 30.75 0.906 0.213 27.89 0.847 0.235 0.453 2.852 0.011